OpenStudy (anonymous):

binomial geometric hypergeometric For each one, Explain the conditions in which we would use it. Justify the formula used. Give an example of a situation in which it could be used

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Hey there ! Please help me!

OpenStudy (anonymous):

Do you know Laino?

OpenStudy (anonymous):

Or it is out of your area of experience?

OpenStudy (michele_laino):

when i was at my university, I used the binomial distribution only

OpenStudy (anonymous):

binominal distribution is used when a system gives equal chance of something to happen right?

OpenStudy (anonymous):

For example the flipping of the coins

OpenStudy (anonymous):

Do you know the formula to binominal distribution?

OpenStudy (michele_laino):

not necessarily. For example, if I have a dice, and i throw that dice say 500 time, then I measure the probability to get the number 3, then the requested probability is: 1/6, whereas the probability to get another number is 5/6

OpenStudy (michele_laino):

so total probability to get a 3 say 200 times, is: \[\Large P\left( 3 \right) = \left( {\begin{array}{*{20}{c}} {500} \\ {200} \end{array}} \right){\left( {\frac{1}{6}} \right)^{200}}{\left( {\frac{5}{6}} \right)^{300}}\]

OpenStudy (michele_laino):

that is a possible application for a binomial distribution

OpenStudy (anonymous):

The formula provided explains what happens when one rolls a dice for 300 times where he gets a 3?

OpenStudy (michele_laino):

that formula, gives us the probability to get the number 3, 200 times, if I throw my dice 500 times

OpenStudy (anonymous):

oh ok. So the formula in itself is already justified.

OpenStudy (anonymous):

You were the brightest pickle in the jar:) Others are just peeking and going away lol

OpenStudy (michele_laino):

thanks! I try my best!!

OpenStudy (anonymous):

binomial -- we have N independent trials with success rate p, and we're interested in the number of successes k out of these N trials; equivalently, we have some objects, and p is the proportion of objects that are special, and we're interested in the number of special objects k we get after picking N such objects if we sample objects *with replacement* (so the proportion p is constant) geometric -- we have N independent trials with success rate p, and we're interested in the number of trials k we need to get *one* success hypergeometric -- we have T objects, and M of them are special, and we're interested in the number of special objects k we get after picking n such objects if we sample *without replacement* (so the proportion changes from M/T with each object we pick)